OpenStudy (anonymous):

The graph of f(x) = 2x + 4 shifts five units to the right when it is replaced with the graph of f(x) = 2x - k. What is the value of k? -5 -1 1 5 @kropot72 @sourwing @sammixboo @iambatman

OpenStudy (anonymous):



OpenStudy (anonymous):

Now if this shifts five to the right, what will your k value become?

OpenStudy (anonymous):

-2 -> 5 right

OpenStudy (anonymous):

-5

OpenStudy (anonymous):

Sec

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

No, wait I think your answer should be y = 2x-6

OpenStudy (anonymous):

But I don't see that option in your solutions mhm

OpenStudy (anonymous):

No that's not it

OpenStudy (anonymous):

it's a +6, the integer is positive because they already told you that function is y = 2x-k so it's a positive integer not negative

OpenStudy (anonymous):

I guess, go with +5

OpenStudy (anonymous):

Idk though because if you graph it out you should get y=2x-6

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Unless I'm wrong or thinking about something else haha, let me tag someone to verify I suppose @ganeshie8 hey could you check this out haha?

ganeshie8 (ganeshie8):

y=2(x-5) +4 = 2x - 10 + 4 = 2x - 6 k = 6

OpenStudy (anonymous):

thats not an option

OpenStudy (anonymous):

I thought so, so the solutions they gave you are wrong, you should ask/ tell your teacher.

OpenStudy (anonymous):

Thanks for checking ganeshie :p

OpenStudy (anonymous):

thanks guys